OpenStudy (anonymous):

can you help me with this solve for the variable and show work ____ V x+6 +2=6

OpenStudy (mdilley):

is your question x+6+2=6?

OpenStudy (anonymous):

no the ___ was suppose to be the symbol for a square root V

OpenStudy (anonymous):

\[\sqrt{x+6} + 2 = 6\] this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do they give you multiple choice answers?

OpenStudy (anonymous):

no were suppose to answer than explain how we got it

OpenStudy (anonymous):

I think I have an idea on what the answer is but it'll suck if they don't give you options haha. Anyways... \[\sqrt{x+6} + 2 = 6\] Subtract 2 from both sides. \[\sqrt{x+6} = 4\] To get rid of the square root, you square it to cancel it out. So, you square it by both sides. \[\sqrt{x+6}^{2} = 4^{2}\] \[x+6 = 16\] The rest should be easy... you solve for x...

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Always welcome! @Redlink12